OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for x in the first equation. 3x + 2y = 7 x - y + 3 = 0 What is the resulting equation? 3x - y - 3 + 2y = 7 3(y - 3) + 2y = 7 3y - 3 + 2y = 7 i got a

OpenStudy (anonymous):

write somthing real quick

OpenStudy (anonymous):

just a letter

OpenStudy (anonymous):

When you times something in a bracket, you need to times the whole thing E1 3x + 2y = 7 E2 x-y+3 = 0 It should be x - y + 3 = 0 x - y = -3 x = y- 3 Sub in to e 1 3(y - 3) + 2y = 7 3y - 9 + 2y = 7 5y = 16 y = 16/5 Sub back into e 2 to find x x - 16/5 + 3 = 0 x - 16/5 = -3 x = -3 + 16/5 x = 1/5 Check using e 2 3(1/5) + 2(16/5) = 7 3/5 + 32/5 = 7 35/5 = 7

OpenStudy (anonymous):

k bye thanks its b i figured it out

OpenStudy (anonymous):

So it is b

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

Sorry....I thought we have to do the whole thing... And then I look at the question again and I am like " oh...... "